# Quantum Sensing Using Cold Atoms and Molecules for Nuclear and Particle Physics

### v2

@@ III.3.2 Varieties of optical lattice clocks
As optical lattice clock technology advances, systems designed for various applications are also developed. These systems extend beyond the conventional setup, where a single atomic ensemble is loaded into a single location in an optical lattice.
@@
Continuous interrogation helps suppress the Dick effect [135, 136], where pulsed interrogation of the atomic system introduces instability at frequencies corresponding to the period of the experiment cycle. In addition, continuous interrogation enables feedback on the phase, causing the Allan deviation to scale rather than . This allows for a faster estimation of systematic shifts. The continuous generation of an ultracold atomic beam has been reported as a step toward developing a continuous optical lattice clock [166].
The rapidly advancing technology of optical tweezer arrays is also applied to Sr optical atomic clocks [167, 168]. A stability study using 88Sr atoms has been reported[101], though characterization of the uncertainties has not been documented. The ability to control individual atoms is advantageous for generating entangled states, which can further enhance stability, as discussed in the next section.
@@ III.3.3 Enhancement of stability by entanglements
The atomic clock system is one of the few systems where entanglement has been demonstrated to enhance their performances. The first proof-of-principle experiments were performed with microwave clocks using 87Rb [169] . Later, larger stability enhancement was achieved with a system where atoms were uniformly coupled to a cavity mode [170]. These studies utilized spin-squeezed states as the source of entanglement. The first spin-squeezed optical clock was demonstrated with Yb[99]. Spin squeezing was performed on the spin-1/2 system in the ground state of 171Yb [171] and then transferred to the excited state of the clock transition to improve clock performance. The study reported a 4.4 dB improvement below the SQL in the stability of the atomic system with a stability of , limited by the frequency fluctuations of the clock laser. An improvement in the clock stability was demonstrated at the stability of [172] with a Sr optical lattice clock. Note that the spin squeezing described here was performed in a cavity quantum electrodynamics (QED) system, where atoms were located in a high-finesse cavity that enhances their coupling to light.
The enhanced stability of optical clocks using Greenberger-Horne-Zeilinger (GHZ) states has been demonstrated with a tweezer array clock.[173] The GHZ state is generated through a strong atom-atom interaction via a Rydberg state. These strongly entangled atoms exhibit high phase sensitivity, leading to improved clock performance when the Ramsey sequence is used. However, because the GHZ and other entangled states are more sensitive to decoherence, atomic coherence time decreases, resulting in shorter interrogation time compared to state-of-the-art optical lattice clocks. This negates the advantage of the reduced QPN. The overall stability of the clock does not necessarily improve with entangled states. Nevertheless, the shorter interrogation time decreases the duration of a single clock cycle, increasing the bandwidth of clock measurements. This increase can be beneficial when the interrogation time is constrained, such as in fountain clocks, or in searches for ultralight dark matter with relatively large masses (see Section VII).
@@ III.4.1 State-of-the-Art Systems
In a conventional ion clock, a single ion is trapped in a Paul trap, which uses an oscillating quadrupole electric field to confine charged particles[174, 175, 176]. Since Paul traps have been in use longer than magic-wavelength optical lattices, ion clocks were developed as the frequency standard before optical lattice clocks, as shown in Fig. 2. Although the 1-s stability for the ion clock is not as high as that of optical lattice clocks, their ability to operate over long periods allows their accuracy to be comparable to that of optical lattice clocks.
The ionic species with the smallest fractional uncertainties have traditionally been Al+ [65] and Yb+ [64]. Lu+ [131], and In+ [79] recently reached the fractional uncertainties of similar level. Al+, In+ and Lu+ have alkaline-earth-atom-like electronic structures, while Yb+ has an alkali-atom-like structure with an E3 transition, where an electron in the orbital is excited. Alkali-atom-like structure with E2 transitions, such as Sr+ [103], Ca+ [102], Ba+ [104], and Ra+ [105] have also been investigated. Some of these clocks are summarized in Table 1. For certain ions, the broad-linewidth transitions suitable for laser cooling have wavelengths too short for readily available laser systems. In such cases, laser cooling is achieved through sympathetic cooling by trapping another ion in the same Paul trap. The preparation and the detection of the quantum state are also carried out using this second ion through the Coulomb coupling of their motional states, which is known as quantum logic spectroscopy [177]. For example, in an Al+ clock, the Al+ ion is sympathetically cooled by a Mg+ ion, and quantum logic spectroscopy is performed via the Mg+ ion [65].
@@ III.4.2 Uncertainties
State-of-the-art ion clocks have fractional uncertainties around , with their accuracy limited by several factors. One of the major sources of systematic uncertainty is the micromotion-related Doppler shift, which arises from the ion’s micromotion in a Paul trap. In some clocks, this uncertainty is the dominant contributor to the overall uncertainties [65, 64, 103]. Suppressing this effect can be achieved by optimizing the Paul trap’s configuration. The high performance of the best Al+ ion clock was made possible by designing a new electrode configuration that minimizes micromotions [65], offering a significant improvement over previous-generation setups.
@@
With respect to the ac Stark shift, the probe-induced ac Stark shift can also be a significant source of uncertainty[64]. This is particularly relevant for ion clocks that use E2 or E3 transitions as the clock transition, as these require much higher laser intensity compared to the , which is weakly mixed with an E1 transition by the nuclear magnetic moment. To mitigate the ac Stark shift induced by the clock laser, the hyper-Ramsey sequence, where pulse lengths are adjusted to cancel the shift, is advantageous[180]. This sequence is particularly beneficial for the spectroscopy requiring high intensity, such as transitions without any mixing with an E1 transition and two-photon transitions. The hyper-Ramsey sequence was experimentally investigated in an Yb+ ion clock using the E3 transition [181]. Even with its application in Ref. [64] the probe-induced ac Stark shift remains a major source of the uncertainty.
@@
Depending on the atomic species and system architecture, other effects can also contribute significantly to uncertainty. The quadratic Zeeman shift is sometimes a major factor[65, 79]. In the case of In+ [79], the uncertainty arises from the relatively large uncertainty in the -factor difference. Al+ has a coefficient for the second-order Zeeman shift more than 10 times larger than that of In+ [182]. The In+ clock identifies thermal time dilation as its major source of uncertainty [79], which can be mitigated by using a narrow-linewidth transition for laser cooling. Other factors, such as background gas collision and servo error, which are common sources of uncertainty in the optical lattice clocks, also contribute to the overall uncertainty.
@@
The main disadvantage of ion clocks is their large QPN due to the small number of ions in the trap, unlike optical lattice clocks which contain a much larger number of neutral atoms. The ultimate solution to reducing QPN is to increase the number of trapped ions. Multi-ion clocks have been investigated both theoretically and experimentally. Theoretically proposal exist for both Paul traps [183] and Penning traps [184]. In Paul traps, a path towards uncertainty below is demonstrated, and the scaling of QPN is reported [79]. A key source of uncertainty in multi-ion clocks is the quadrupole shift, which arises because ions are not located at the center of the Paul trap. As a result, each ion experiences a different quadrupole electric field, leading to inhomogeneous broadening of the transition. This effect is particularly problematic for ions with large quadrupole moments, such as Ca+, Sr+, and Ba+ [185]. However, Al+ and In+ are much less affected due to their small quadrupole moments [186]. Suppression of this quadrupole shift has been demonstrated using dynamic decoupling [187]. Also, quantum logic spectroscopy has been successfully applied to multiple ions[188].
@@ IV.1 Theoretical basis
Searches for variations of fundamental constants have a long history. Early investigations were performed through astronomical observations [197] and studies of the natural nuclear reactor at Oklo[198]. Since than, extensive astronomical searches for variations of fundamental constants have been long ongoing[196]. Particularly important were observations based on atomic absorption lines which suggested that may have had a different value in the region of universe with large redshift . A study in 2001 [199] reported over , where is the variation of , wit no identified systematic effects that could account for the results. Subsequent analyses incorporating additional data have supported spatial variation of with significance (see Ref.[200] and references therein). However, recent observations at large themselves are consistent with no variation [200, 201, 202]. If this spatial variation extends to the scale of the solar system, a variation of yr-1 is expected due to the motion of the solar system in the universe [203]. Comparisons using atomic clocks provide a valuable cross-check against astronomical searches, as they allow for better control of environment of atoms than astronomical searches.
@@ IV.2 Status of experimental searches in laboratories
To increase the sensitivity to the time variation of in a fixed amount of time, either the stability of an atomic clock or has to be enhanced. Notably, the state in neutral Yb can increase sensitivity [117, 118, 208]. Although for this transition is not the largest in Table 2, optical lattice clocks have higher short-term stability than ion clocks, which makes this state beneficial for enhancing sensitivity, especially for short-term variations. Note that the large ’s in the and transitions are mainly due to their small transition energy, and the excited state is the same as the transition. In this case, the overall sensitivity for the variation of needs to be carefully estimated for unstable clocks. Ref. [117] argues that the enhancement in is cancelled by the clock instability, if the instability is mainly due to noise on the . To take advantage of transitions with large , transitions in highly charged ions and 229Th are promising.
@@ V.1 Properties of highly charged ions
A major source of systematic uncertainty in optical clocks arises from the ac Stark shift, which is caused by the deformation of the electronic cloud around the nucleus due to an external electric field. This effect can be mitigated by reducing the deformation, which is achieved by holding electrons more tightly within the atom. Such strong confinement of electrons can be achieved in a highly charged ion (HCI) [232]. Qualitatively, by stripping away loosely bound the outer-shell electrons, only those with strong binding remain, leading to reduced susceptibility to external electric fields. Optical clocks based on HCIs have been proposed as a next-generation clock capable of surpassing the accuracy limit of both optical lattice clocks and ion clocks [17].
@@
A key tradeoff of HCI clocks compared to optical lattice clocks is the number of ions that can be interrogated simultaneously. HCI clocks typically trap a single HCI,resulting in lower clock stability due to QPN. A HCI is more susceptible to deionization from collisions with residual gas than a singly charged ion, which potentially can limit the duration of long-term operation.
@@
Beyond their role in next-generation atomic clocks, HCIs are also useful for various fundamental physics searches. One example is testing QED in strong fields. The intense electric field experienced by electrons in U90+ is used to probe QED in extreme conditions[234]. In addition, HCIs exhibit higher sensitivity to the variation of than neutral atoms. This is because high-velocity electrons in HCIs are affected more by relativistic effects, which scale strongly with . Following a pioneering work [235],various ionic species have been proposed as suitable candidates for searching for the variation of . Several transitions with large are listed in Table 2, with additional transitions provided in Table 5. Transitions with longer wavelengths tend to have large , as explained by Eq. 6. Another important application is the search for fifth force between an electron and a neutron (See Sec. VIII for further details). Due to their significantly different electronic structures compared to neutral atoms, HCIs offer high sensitivity for detecting new forces through isotope shift measurements [236].
@@
To perform laser spectroscopy on HCIs, both the atomics species and must be properly selected. The scaling of binding energy suggests that electronic transitions between different states, which typically occur in the visible-light range for neutral atoms, shift into the X-ray region for HCIs. Instead, energy differences between hyperfine or fine structures increase to a range suitable for laser spectroscopy. For hyperfine structures, hydrogen-like ions of heavy atoms exhibit eV energy differences. Table. 4 summarizes ions with nuclear spin and transition frequencies accessible by lasers. For fine structures, eV energy differences are realized at . An example is the energy-level structure of Pr10+, as shown in Fig. 5, where the 255.6 nm transition is between two fine-structure levels. The two states are another pair of fine structures. The transition between J=3 and J=2 states in the state of Cf16+, listed in Table 5, is also a transition between fine structures.
Naively, energy differences between different orbitals exceed the energy for visible light. However, an exceptional case where visible light can drive electronic transitions in HCIs is known as level crossing [17]. Fig. 6 provides an intuitive illustration of this phenomenon. In a hydrogen atom, different states for the same level are degenerate, where is the azimuthal quantum number. However, in multielectron atoms, the states are ordered from the low energy as . Consider, for example, the energy of the and orbitals. In a hydrogen-like ion, the orbital has a higher energy than the orbital, whereas in a neutral atom, the orbital has higher energy. At a certain in HCIs, the energy difference between these two states reverses sign. Around this , the energy of the two orbitals is nearly the same, as shown in Fig. 6, sometimes falling into the visible light range.

### v3

@@ III.3.2 Varieties of optical lattice clocks
As the optical lattice clock technology advances, systems designed for various applications are also developed. These systems extend beyond the conventional setup, where a single atomic ensemble is loaded into a single location in an optical lattice.
@@
Continuous interrogation helps suppress the Dick effect [138, 139], where pulsed interrogation of the atomic system introduces instability at frequencies corresponding to the period of the experiment cycle. In addition, continuous interrogation enables feedback on the phase, causing the Allan deviation to scale rather than . This allows for a faster estimation of systematic shifts. A continuous generation of an ultracold atomic beam has been reported as a step toward developing a continuous optical lattice clock [172].
The rapidly advancing technology of optical tweezer arrays is also applied to Sr optical atomic clocks [173, 174]. A stability study using 88Sr atoms has been reported[175], though characterization of the uncertainties has not been documented. The ability to control individual atoms is advantageous for generating entangled states, which can further enhance stability, as discussed in Sec. III.3.3.
@@ III.3.3 Enhancement of stability by entanglements
The atomic clock system is one of the few systems where entanglement has been demonstrated to enhance their performance. The first proof-of-principle experiments were performed with microwave clocks using 87Rb [176], where atoms trapped in an optical cavity are entangled through a strong interaction with photons circulating in the optical cavity. Later, larger stability enhancement was achieved with a system where atoms were uniformly coupled to a light field inside an optical cavity [177]. These studies utilized spin-squeezed states as the source of entanglement. The first spin-squeezed optical atomic clock was demonstrated with Yb[102]. Spin squeezing was performed on the spin-1/2 system in the ground state of 171Yb [178] and then transferred to the excited state of the clock transition to improve clock performance. The study reported a 4.4 dB improvement below the SQL in the stability of the atomic system with a fractional stability of , limited by the frequency fluctuations of the clock laser. An improvement in the clock stability was demonstrated at the stability of [179] with a Sr optical lattice clock. With improved stability of the system, longer interrogation time and improved metrological gain is achieved in the same system. This resulted in the spin squeezed optical lattice clock beyond the SQL at the fractional stability of [180]. Note that all spin squeezing described here was performed in a cavity quantum electrodynamics (QED) system, where atoms were located in a high-finesse cavity that enhances their coupling to light.
Enhanced stability of optical atomic clocks using Greenberger-Horne-Zeilinger (GHZ) states has been demonstrated with a tweezer array clock.[181] The GHZ state is generated through a strong atom-atom interaction via a Rydberg state. These strongly entangled atoms exhibit high phase sensitivity, leading to improved clock performance when the Ramsey sequence is used. However, because the GHZ and other entangled states are more sensitive to decoherence, the atomic coherence time decreases, resulting in shorter interrogation time compared to state-of-the-art optical lattice clocks. This negates the advantage of the reduced QPN. The overall stability of the clock does not necessarily improve with entangled states. Nevertheless, the shorter interrogation time decreases the duration of a single clock cycle, increasing the bandwidth of clock measurements. This increase can be beneficial when the interrogation time is constrained, such as in fountain clocks, or in searches for ultralight dark matter with relatively large masses (see Section VII).
@@ III.4.1 State-of-the-Art Systems
In a conventional ion clock, a single ion is trapped in a Paul trap, which uses an oscillating quadrupole electric field to confine charged particles[182, 183, 184]. Since Paul traps have been in use for a longer time than magic-wavelength optical lattices, ion clocks were developed as the frequency standard before optical lattice clocks, as shown in Fig. 2. Although the one-second stability of the ion clock is not as high as that of optical lattice clocks, their ability to operate over long periods allows their accuracy to be comparable to that of optical lattice clocks.
The ionic species with the smallest fractional uncertainties have traditionally been Al+ [65] and Yb+ [64]. Lu+ [133], In+ [82], Ca+[69], and Sr+[68] recently reached the fractional uncertainties of a similar level. Al+, In+, and Lu+ have alkaline-earth-atom-like electronic structures, while Yb+, Ca+, and Sr+ have alkali-atom-like structures. Yb+ clocks make use of the E3 transition, where an electron in the orbital is excited. Sr+ and Ca+ utilize E2 transitions. Other atoms with alkaline-atom-like structure such as Ba+ [106] and Ra+ [107] have also been investigated. Some of these clocks are summarized in Table 1. For certain ions, the broad-linewidth transitions suitable for laser cooling have wavelengths too short for readily available laser systems. In such cases, laser cooling is achieved through sympathetic cooling by trapping another ion in the same Paul trap. The preparation and the detection of the quantum state are also carried out using this second ion through the Coulomb coupling of their motional states, which is known as quantum logic spectroscopy [185]. For example, in an Al+ clock, the Al+ ion is sympathetically cooled by a Mg+ ion, and quantum logic spectroscopy is performed via the Mg+ ion [65].
@@ III.4.2 Uncertainties
State-of-the-art ion clocks have fractional uncertainties around . An Yb+ clock first reached in 2016[64], and an Al+ clock reached in 2019. In 2025, an In+[82] with is reported. Also, three ion clocks achieved fractional uncertainties below : for Al+[67], for Sr+[68], and for Ca+[69]. Accuracy of these clocks is limited by several factors. One of the major sources of systematic uncertainty is the micromotion-related Doppler shift, which arises from the ion’s micromotion in a Paul trap. In some clocks, this uncertainty is the dominant contributor to the overall uncertainties [65, 64, 105]. Suppressing this effect can be achieved by optimizing the Paul trap’s configuration. The high performance of the best Al+ ion clock was made possible by designing a new electrode configuration that minimizes micromotions [65, 67], offering a significant improvement over previous-generation setups.
@@
With respect to the ac Stark shift, the probe-induced ac Stark shift can also be a significant source of uncertainty[64]. This is particularly relevant for ion clocks that use E2 or E3 transitions as the clock transition, as these require much higher laser intensity compared to the transitions, which are weakly mixed with an E1 transition by the nuclear magnetic moment. To mitigate the ac Stark shift induced by the clock laser, the hyper-Ramsey sequence, where pulse lengths are adjusted to cancel the shift, is advantageous[188]. This sequence is particularly beneficial for the spectroscopy requiring high intensity, such as transitions without any mixing with an E1 transition and two-photon transitions. The hyper-Ramsey sequence was experimentally investigated in an Yb+ ion clock using the E3 transition [189]. Even with its application in Ref. [64] the probe-induced ac Stark shift remains a major source of the uncertainty.
@@
Depending on the atomic species and system architecture, other effects can also contribute significantly to uncertainty. The quadratic Zeeman shift is sometimes a major factor[65, 82]. Al+ has a coefficient for the second-order Zeeman shift more than 10 times larger than that of In+ [190]. In the case of In+ [82], the uncertainty arises from the relatively large uncertainty in the -factor difference. In addition, the In+ clock also identifies thermal time dilation as its major source of uncertainty [82], which can be mitigated by using a narrow-linewidth transition for laser cooling. Other factors, such as background gas collision and servo error, which are common sources of uncertainty in the optical lattice clocks, also contribute to the overall uncertainty.
@@
The main disadvantage of ion clocks is their large QPN due to the small number of ions in the trap, unlike optical lattice clocks which contain a much larger number of neutral atoms. The ultimate solution to reducing QPN is to increase the number of trapped ions. Multi-ion clocks have been investigated both theoretically and experimentally. Theoretically proposal exist for both Paul traps [191] and Penning traps [192]. In Paul traps, a path towards uncertainty below is demonstrated, and the scaling of QPN is reported [82]. A key source of uncertainty in multi-ion clocks is the quadrupole shift, which arises from electric field gradients due to trap electrodes and Coulomb interaction between ions. As a result, each ion experiences a different quadrupole electric field, leading to inhomogeneous broadening of the transition. This effect is particularly problematic for ions with large quadrupole moments, such as Ca+, Sr+, and Ba+ [193]. However, Al+ and In+ are much less affected due to their small quadrupole moments [194]. Suppression of this quadrupole shift has been demonstrated using dynamic decoupling [195]. Also, quantum logic spectroscopy has been successfully applied to multiple ions[196].
With an advent of multi-ion clocks, enhancing the stability of an ion clock by entanglement is attempted [197]. Because it is easier to achieve maximally entangled states with few ions than many atoms (See Fig. 2 of Ref. [198]; entanglement at the Heisenberg limit is achieved by trapped ions), and because the leading factor limiting the QPN of ion clocks is the number of ions, enhancement of the phase sensitivity by the entanglement is effective to improve the stability of ion clocks. In Ref. [197], a two-ion Ca+ clock is operated with entanglement, and the highest one-second stability is achieved for Ca+ clocks to date.
@@ IV.1 Theoretical basis
Searches for variations of fundamental constants have a long history. Early investigations were performed through astronomical observations [207] and studies of the natural nuclear reactor at Oklo[208]. Since than, extensive astronomical searches for variations of fundamental constants have been long ongoing[206, 209]. Particularly important were observations based on atomic absorption lines that suggested that may have had a different value in the region of universe with a large redshift . A study in 1999 [210] claimed over . A subsequent report in 2001 [211] showed an increased significance with over , where is the variation of , with no identified systematic effects that could account for the results. Further analyses incorporating additional data have supported spatial variation of with significance (see Ref.[212] and references therein). However, recent observations at large themselves are consistent with no variation [212, 213, 214]. If this spatial variation extends to the scale of the solar system, a variation of yr-1 is expected due to the motion of the solar system in the universe [215]. Comparisons using atomic clocks provide a valuable cross-check against astronomical searches, as they allow for better control of environment of atoms than astronomical searches.
@@ IV.2 Status of experimental searches in laboratories
To increase the sensitivity to the time variation of in a fixed amount of time, either the stability of an atomic clock or has to be enhanced. Notably, the state in neutral Yb can increase sensitivity [120, 121, 220]. Although for the transition from the ground state to this state is not the largest in Table 2, optical lattice clocks have higher short-term stability than ion clocks, which makes this state beneficial for enhancing sensitivity, especially for short-term variations. Note that the large ’s in the and transitions are mainly due to their small transition energy, and the excited state is the same as the transition. In this case, the overall sensitivity for the variation of needs to be carefully estimated for unstable clocks. Ref. [120] argues that the enhancement in is cancelled by the clock instability, if the instability is mainly due to noise on the . To take advantage of transitions with large , transitions in highly charged ions and 229Th are promising.
@@ V.1 Properties of highly charged ions
A major source of systematic uncertainty in optical atomic clocks arises from the ac Stark shift, which is caused by the deformation of the electronic cloud around the nucleus due to external electric fields. This effect can be mitigated by reducing the deformation, which can be achieved by holding electrons more tightly within the atom. Such strong confinement of electrons can be achieved in a highly charged ion (HCI) [245]. Qualitatively, by stripping away loosely bound outer-shell electrons, only those with strong binding remain, leading to reduced susceptibility to external electric fields. Optical atomic clocks based on HCIs have been proposed as a next-generation clock capable of surpassing the accuracy limit of both optical lattice clocks and ion clocks [17].
@@
A key trade-off of HCI clocks compared to optical lattice clocks is the number of ions that can be interrogated simultaneously. HCI clocks typically trap a single HCI, resulting in lower clock stability due to QPN. A HCI is more susceptible to deionization from collisions with residual gas than a singly charged ion, which potentially can limit the duration of long-term operation.
@@
Beyond their role in next-generation atomic clocks, HCIs are also useful for various fundamental physics searches. One example is testing QED in strong fields. The intense electric field experienced by electrons in innermost orbitals is suitable for probing QED in strong-field regime. Recent reports include spectroscopy of helium-like U90+ [247, 248] and beryllium-like Pb78+ [249]. In addition, HCIs exhibit higher sensitivity to the variation of than neutral atoms. This is because high-velocity electrons in HCIs are affected more by relativistic effects, which scale strongly with . Following a pioneering work [250],various ionic species have been proposed as suitable candidates for searching for the variation of . Several transitions with large are listed in Table 2, with additional transitions provided in Table 5. Transitions with longer wavelengths tend to have large , as explained by Eq. 5. Another important application is the search for fifth force between an electron and a neutron (See Sec. VIII for further details). Due to their significantly different electronic structures compared to neutral atoms, HCIs offer high sensitivity for detecting new forces through isotope shift measurements [251].
@@
To perform laser spectroscopy on HCIs, both the atomic species and must be properly selected. The scaling of binding energy suggests that electronic transitions between different states, which typically occur in the visible-light range for neutral atoms, shift into the x-ray region for HCIs. Instead, energy differences between hyperfine or fine structures increase to a range suitable for laser spectroscopy. For hyperfine structures, hydrogen-like ions of heavy atoms exhibit eV energy differences. Table. 4 summarizes ions with nuclear spin and transition frequencies accessible by lasers. For fine structures, eV energy differences are realized at . An example is the energy-level structure of Pr10+, as shown in Fig. 5, where the 255.6 nm transition is between two fine-structure levels. The two states are another pair of fine structures. The transition between J=3 and J=2 states in the state of Cf16+, listed in Table 5, is also a transition between fine structures.
Naively, energy differences between different orbitals exceed the energy for visible light. However, an exceptional case where visible light can drive electronic transitions in HCIs is known as level crossing [17]. Figure 6 provides an intuitive illustration of this phenomenon. In a hydrogen atom, different states for the same level are degenerate, where is the azimuthal quantum number. However, in multielectron atoms, the states are ordered from the low energy as . Consider, for example, the energy of the and orbitals. In a hydrogen-like ion, the orbital has a higher energy than the orbital, whereas in a neutral atom, the orbital has higher energy. At a certain in HCIs, the energy difference between these two states reverses sign. Around this , the energy of the two orbitals is nearly the same, as shown in Fig. 6, sometimes falling into the visible light range.